OpenStudy (anonymous):

Finding complex solutions for 9x^2 - 15x + 25 = 0 How do I obtain the simplest complex solutions from this?

OpenStudy (anonymous):

use the quadratic formula \[(-b \pm \sqrt{b^{2}-4ac)}/2a\] in this case a=9, b=-15 and c= 25

OpenStudy (anonymous):

your should get something like \[\frac{ 15\pm \sqrt[2]{-675} }{ 50 }\]

OpenStudy (anonymous):

and then use i for the negative sqr root\[\frac{ 15 }{ 50 }\pm \frac{ \sqrt{675\times \sqrt{-1}} }{ 50 }=\] \[.3\pm.5196152423i\]

OpenStudy (anonymous):

Thanks! I forgot about the quadratic formula lol.